OpenStudy (anonymous):

the vertecies of a quadilateral are: (2,2), (4,6), (8,2) and (10,6). which BEST describes this quadilateral? A. Parallelogram B. Recatangle C. Rhombus D. Square

OpenStudy (ranga):

You can graph the points on a graph paper and figure it out. But if you want to figure it out analytically then: Call the vertices ABCD in order. First step: Find the lengths: AB, BC, CD and DA.

OpenStudy (anonymous):

thank you sooo much!!!! so i found the distance between each set of the pints so now what do i do?? @ranga

OpenStudy (ranga):

What are the distances? AB, BC, CD and DA?

OpenStudy (anonymous):

AB= 4.47 BC= 5.65 CD= 4.47 DA= 8.9 im not sure if any of them are wrong. @ranga

OpenStudy (ranga):

A: (2,2), B:(4,6), C: (8,2) and D: (10,6) Just realized they have tricked us by switching the naming order of the vertices. If we look at A and C, they have the same y coordinate of 2. That means A and C are on the same horizontal line. If we look at B and D, they have the same y coordinate of 6. That means B and D are on the same horizontal line. So we have a 4 sided figure AC at the bottom and BD at the top and they are both horizontal lines and therefore they are parallel. So the order of vertices is: ACDB. So we need to find the distances AC, CD, DB and BA to find the length of the sides.

OpenStudy (ranga):

The quadrilateral is ACDB A: (2,2), B:(4,6), C: (8,2) and D: (10,6) AC = 8 - 2 = 6 CD = sqrt{ (10-8)^2 + (6-2)^2 } = sqrt(4+16) = sqrt(20) DB =10 - 4 = 6 BA = sqrt{ (6-2)^2 + (4-2)^2 } = sqrt(16+4) = sqrt(20) The 4 sides are not of equal length and that rules out square and rhombus.

OpenStudy (ranga):

We know AC and BD are horizontal lines and their slope is 0. Let us find the slope of the line AB. A: (2,2), B:(4,6) slope of AB = (6-2)/(4-2) = 4/2 = 2 The side AB has a slope of 2. Its is NOT a vertical side. Therefore it is NOT a rectangle. The given points form a parallelogram. Their opposite sides are equal and parallel.

OpenStudy (anonymous):

okay so i got AC=6 CD=4.47 DB=6 BA=4.47 so that makes alllott of sence!! Thank You sooo Much!!!! ^^ @ranga

OpenStudy (ranga):

You are welcome. For these kinds of problems, plotting it in a graph paper or a graphing calculator or even an online graphing website such as https://www.desmos.com/calculator may be the quickest way to figure out what shape the points represent.

OpenStudy (anonymous):

ooh okay!!! Thank you thaat would be much more useful next time @ranga

OpenStudy (ranga):

Alright. You are welcome.